OpenStudy (anonymous):

Medal award! Find the arc length along a circle of diameter 9.5 meters which subtends an angle of 1.9 radians

OpenStudy (anonymous):

diameter=9.5 m radius r=?

OpenStudy (anonymous):

4.75

OpenStudy (anonymous):

correct. \[\theta=\frac{ l }{ r },l=r \times \theta \] where l= length of arc r=radius if circle \[\theta=angle~of~arc(in~radians)\]

OpenStudy (anonymous):

can you solve?

OpenStudy (anonymous):

so I have 4.75*9.5?

OpenStudy (anonymous):

4.75*1.9

OpenStudy (anonymous):

correct.

OpenStudy (anonymous):

so I would just need to input 1.9pi*4.75? Because I tried that and it mentioned that it was incorrect

OpenStudy (anonymous):

4.75*1.9=9.025 m

OpenStudy (anonymous):

thanks so much :)

OpenStudy (anonymous):

why are you writing pi ? angle is already given in radians.

OpenStudy (anonymous):

when angle is given in degrees then you need to change it in radians.

OpenStudy (anonymous):

yep :) I messed up but the 9.025 is correct! Thanks so much

OpenStudy (anonymous):

yw